### v1

# StylizedGS: Controllable Stylization for 3D Gaussian Splatting

Abstract
With the rapid development of XR, 3D generation and editing are becoming more and more important, among which, stylization is an important tool of 3D appearance editing. It can achieve consistent 3D artistic stylization given a single reference style image and thus is a user-friendly editing way. However, recent NeRF-based 3D stylization methods face efficiency issues that affect the actual user experience and the implicit nature limits its ability to transfer the geometric pattern styles. Additionally, the ability for artists to exert flexible control over stylized scenes is considered highly desirable, fostering an environment conducive to creative exploration. In this paper, we introduce StylizedGS, a 3D neural style transfer framework with adaptable control over perceptual factors based on 3D Gaussian Splatting (3DGS) representation. The 3DGS brings the benefits of high efficiency. We propose a GS filter to eliminate floaters in the reconstruction which affects the stylization effects before stylization. Then the nearest neighbor-based style loss is introduced to achieve stylization by fine-tuning the geometry and color parameters of 3DGS, while a depth preservation loss with other regularizations is proposed to prevent the tampering of geometry content. Moreover, facilitated by specially designed losses, StylizedGS enables users to control color, stylized scale and regions during the stylization to possess customized capabilities. Our method can attain high-quality stylization results characterized by faithful brushstrokes and geometric consistency with flexible controls. Extensive experiments across various scenes and styles demonstrate the effectiveness and efficiency of our method concerning both stylization quality and inference FPS.
Index Terms:
Gaussian Splatting, Style Transfer, Perceptual Control
Figure 1: Stylization Results. Given a 2D style image, the proposed StylizedGS method can stylize the pre-trained 3D Gaussian Splatting to match the desired style with detailed geometric features and satisfactory visual quality within a few minutes. We also enable users to control several perceptual factors, such as color, the style pattern size (scale), and the stylized regions (spatial), during the stylization to enhance the customization capabilities.
1 Introduction
Nowadays, the once professionally-dominated domain of artistic content creation has become increasingly accessible to novice users, thanks to recent groundbreaking advancements in visual artistic stylization research. As a pivotal artistic content generation tool in crafting visually engaging and memorable experiences, 3D scene stylization has attracted growing research efforts. Previous methodologies have attempted style transfer by enabling control over diverse explicit representations such as mesh [1, 2, 3], voxel [4, 5], and point cloud [6, 7, 8]. However, the quality of their results is limited by the quality of the geometric reconstructions. The recent 3D stylization methods benefit from the emerging implicit neural representations [9, 10, 11], such as neural radiance field (NeRF) [12, 13, 14, 15, 16], achieving more faithful and consistent stylization within 3D scenes. Nonetheless, NeRF-based methods are computationally intensive to optimize and suffer from the geometry artifacts in the original radiance fields.
The recently introduced 3D Gaussian Splatting (3DGS) [17], showcasing remarkable 3D reconstruction quality from multi-view images with high efficiency, suggests representing the 3D scene using an array of colored and explicit 3D Gaussians. Given that practical 3D stylization applications often demand a prompt response, we propose using 3DGS as the representation of real-world scenes and conducting 3D stylization on it. Recent 3DGS scene manipulation methods [18, 19, 20, 21] explore the editing and control of 3D Gaussians using text instructions within designated regions of interest or semantic tracing. However, these approaches are constrained by text input and fall short of delivering detailed style transfer capabilities. The methods for performing 3D Gaussian stylization remain less explored in the field.
In this paper, we introduce the first controllable scene stylization method based on 3DGS, StylizedGS. With a single reference style image, our method can effectively transfer its style features to the entire 3D scene, represented by a set of 3D Gaussians. It facilitates the artistic creation of visually coherent novel views that exhibit transformed and detailed style features in a visually reasonable manner. More importantly, StylizedGS operates at a notable inference speed, ensuring the efficient synthesis of stylized scenes. To address the desire for users to control perceptual factors such as color, scale, and spatial aspects, as in 2D image style transfer [22, 23, 24] and 3D scene style transfer [25], we enhance the flexibility and introduce an advanced level of perceptual controllability in our approach to achieve personalized and diverse characteristics in 3DGS stylization.
The proposed 3DGS stylization process is formulated as an optimization problem that optimizes the geometry and color of 3D Gaussians to render images with faithful stylistic features while preserving the semantic content. A trivial solution is directly minimizing both a style loss and a content loss to align the rendered images with the style image and avoid overly strong stylization by fine-tuning the color of 3D Gaussians, respectively. Nevertheless, it poses a challenge in obtaining high-quality results with intricate style details as shown in Fig. 10 (b), as learning only the color alterations cannot effectively capture the overall style pattern. Instead, we propose to learn both the optimal geometry and color parameters of 3D Gaussians to capture the detailed style feature and facilitate the stylization of the entire 3D scene. We propose a 2-step stylization framework with the first step to enhance the fidelity in the stylized scene by the color match and reduce cloudy artifacts and geometry noises by the proposed GS filter. The second step is the optimization for stylization with a nearest neighbor feature match (NNFM) loss. We also introduce a depth preservation loss without the need for additional networks or regulation operations to preserve the overall learned 3D scene geometry. Last, we specially design a set of effective loss functions and optimization schemes to enable flexible perceptual control for users, including color, scale, and spatial regions Our contribution can be summarized as follows:
  • We introduce StylizedGS, a novel controllable 3D Gaussian stylization method that organically integrates various modules with proper improvements to transfer detailed style features and produce faithful novel stylized views.
  • We empower users with an efficient stylization process and flexible control by specially designed losses, enhancing their creative capabilities.
  • Our approach achieves significantly reduced training and rendering times while generating high-quality stylized scenes compared with existing 3D stylization methods.
2 Related Work
Image Style Transfer. Style transfer aims to generate synthetic images with the artistic style of given images while preserving content. Initially proposed in neural style transfer methods by Gatys et al. [26, 27], this process involves iteratively optimizing the output image using Gram matrix loss and content loss calculated from VGG-Net [28] extracted features. Subsequent works [29, 30, 31, 32] have explored alternative style loss formulations to enhance semantic consistency and capture high-frequency style details such as brushstrokes. Feed-forward transfer methods [33, 34, 35], where neural networks are trained to capture style information from the style image and transfer it to the input image in a single forward pass, ensuring faster stylization. Recent improvements in style loss [31, 32, 13] involve replacing the global Gram matrix with the nearest neighbor feature matrix, improving texture preservation. Some methods adopt patch matching for image generation, like Fast PatchMatch [36] and PatchMatch [37], but are limited to specific views. Combining neural style transfer with novel view synthesis methods without considering 3D geometry may lead to issues like blurriness or view inconsistencies.
3D Gaussian Splatting. 3D Gaussian Splatting (3DGS) [17] has emerged as a robust approach for real-time radiance field rendering and 3D scene reconstruction. Recent methods [20, 19, 18, 38] enhance semantic understanding of 3D scenes and enabled efficient text-based editing using pre-trained 2D models. Despite these advancements, existing 3DGS works lack support for image-based 3D scene stylization that faithfully transfers detailed style features while offering flexible control. In this work, leveraging the 3DGS representation, we significantly reduce both training and rendering times for stylization, enabling interactive perceptual control over stylized scenes.
Figure 2: StylizedGS Pipeline. We first reconstruct a photo-realistic 3DGS $G_{\theta}^{rec}$ from multi-view input. Following this, color matching with the style image is performed, accompanied by the application of 3DGS filters to preemptively address potential artifacts. Throughout optimization, we employ multiple loss terms to capture detailed local style structures and preserve geometric attributes. Users can flexibly control color, scale, and spatial attributes during stylization through customizable loss terms. Once this stylization is done, we can obtain consistent free-viewpoint stylized renderings.
3D Style Transfer. 3D scene style transfer aims to transfer the style to the scene with both style fidelity and multi-view consistency. With the increasing demand for 3D content, neural style transfer has been expanded to various 3D representations. Stylization on meshes often utilizes differential rendering to propagate style transfer objectives from rendered images to 3D meshes, enabling geometric or texture transfer [3, 1, 2]. Other works, using point clouds as the 3D proxy, ensure 3D consistency when stylizing novel views. For instance, [7] employs featurized 3D point clouds modulated with the style image, followed by a 2D CNN renderer to generate stylized renderings. However, explicit methods’ performance is constrained by the quality of geometric reconstructions, often leading to noticeable artifacts in complex real-world scenes.
Therefore implicit methods such as NeRF [39] have gained considerable attention for their enhanced capacity to represent complex scenes. Numerous NeRF-based stylization networks incorporate image style transfer losses [26, 13] during training or are supervised a mutually learned image stylization network [12] to optimize color-related parameters based on a reference style. Approaches like [14, 40] support both appearance and geometric stylization to mimic the reference style, achieving consistent results in novel-view stylization. However, these methods involve time-consuming optimization and exhibit slow rendering due to expensive random sampling in volume rendering. They also lack user-level flexible and accurate perceptual control for stylization.
While controlling perceptual factors, such as color, stroke size, and spatial aspects, has been extensively explored in image domain style transfer [22, 23, 24], the application of perceptual control in 3D stylization has not been well utilized. [41, 42] establish semantic correspondence in transferring style across the entire stylized scene but only limited to spatial control and doesn’t allow users to interactively specify arbitrary regions. ARF-plus [25] introduce more perceptual controllability into the stylization of radiance fields, yet the demand for enhanced flexibility and personalized, diverse characteristics in 3D stylization remains unmet.
Our approach achieves rapid stylization within a minute of training, ensuring real-time rendering capabilities. It adeptly captures distinctive details from the style image and preserve recognizable scene content with fidelity. Additionally, we empower users with perceptual control over color, scale, and spatial factors for customized stylization.
3 Method
Given a 3D scene represented by a collection of captured images with corresponding camera parameters, our objective is to efficiently accomplish consistent style transfer from a 2D style image with arbitrary style to the 3D scene. Building on recent advances in 3D Gaussian Splatting (3DGS) [17], we model the 3D scene with the 3DGS and then optimize the 3DGS representation to generate stylized scenes with intricate artistic features and high visual quality. Our stylization process consists of two key steps: 1) We first recolor the 3D scene to align its color statistics with those of the given style image. Simultaneously, a specific 3D Gaussian filter is applied to minimize the impact of floaters in the reconstruction, which is crucial for the final stylization effect. 2) Then in the optimization phase, we exploit nearest-neighbor feature matching style loss to capture detailed local style patterns. Additionally, we incorporate a depth preservation loss, and some regularization terms to preserve geometric contents. Finally, in Sec. 3.3, we introduce how to achieve flexible control over the color, scale, and spatial attributes in our 3DGS stylization. As a result, users can create stylized renderings with customized artistic expression, and explore the consistently stylized scene in a free-view manner.
3.1 Preliminaries: 3D Gaussian Splatting
Gaussian Splatting [17] encapsulates 3D scene information using a suite of 3D colored Gaussians. This technique exhibits rapid inference speeds and exceptional reconstruction quality compared to NeRF. To represent the scene, each Gaussian is described by a centroid $p=\{x,y,z\}\in\mathbb{R}^{3}$, a 3D vector $s\in\mathbb{R}^{3}$ for scaling, and a quaternion $q\in\mathbb{R}^{4}$ for rotation. Additionally, an opacity value $\alpha\in\mathbb{R}$ and a color vector $c$ represented in the coefficients of a spherical harmonic (SH) function of degree three are used for fast alpha-blending during rendering. These trainable parameters are collectively symbolized by $G_{\theta_{i}}$, where $G_{\theta_{i}}={p_{i},s_{i},q_{i},\alpha_{i},c_{i}}$, representing the parameters for the $i$-th Gaussian. To visualize the 3D Gaussians and supervise their optimization, 3DGS projects them onto the 2D image plane. The implementation leverages differentiable rendering and gradient-based optimization on each pixel for the involved Gaussians. The pixel color $c^{\alpha}$ is determined by blending the colors $c_{i}$ of those ordered Gaussians that overlap the pixel. This process can be formulated as:
$c^{\alpha}=\sum_{i\in N}T_{i}\alpha_{i}c_{i}$ (1)
where $T_{i}$ is the accumulated transmittance and $\alpha_{i}$ is the alpha-compositing weight for the $i$-th Gaussian.
3.2 Style Transfer to 3D Gaussian Splatting
Given a set of multi-view images $\mathcal{I}_{content}$, we first obtain the reconstructed 3DGS model $G_{\theta}^{rec}$ to represent the original scene. Our goal is to transform $G_{\theta}^{rec}$ to a stylized 3DGS model $G_{\theta}^{sty}$ which matches the detailed style features of a 2D style image $\mathcal{I}_{style}$ and preserve the content of the original scene.
Color Match. To enhance the alignment of hues with the style image and achieve flexible color control in the stylized results, we initially recolor the 3DGS $G_{\theta}^{rec}$ through color transfer from the style image. Drawing inspiration from [22], we employ a linear transformation of colors in RGB space, followed by a color histogram matching procedure between the style images and training views. Let $\{p_{s}^{i}\}$ be the set of all pixels in the style image, $\{p_{c}^{i}\}$ be the set of all pixels in the content image to be recolored, then we solve the following linear transformation to align the mean and covariance of color distributions between the content image set and the style image:
$p_{c}^{re}=\mathbf{A}p_{c}+b,\,c^{re}=\mathbf{A}c+b$
$s.t.\,E[p_{c}^{re}]=E[p_{s}],Cov[p_{c}^{re}]=Cov[p_{s}]$
where $\mathbf{A}\in\mathbb{R}^{3\times 3}$ and $b\in\mathbb{R}^{3}$ are the solution to the linear transformation, $p_{c}^{re}$ is the the recolored pixels of the content images. The color parameter $c$ of 3DGS $G_{\theta}^{rec}$ is transformed to the recolored color parameter $c^{re}$. Please refer to the supplementary document for the mathematical derivation.
Consequently, this alignment process ensures that the recolored training images $\mathcal{I}_{content}^{re}$ and the color attributes of 3DGS $c^{re}$, are consistent with the color palette of the style reference. However, some floaters in the original 3DGS will also be colored during this process, significantly affecting the quality of the stylization, as shown in Fig. 11. These colored floaters often result from Gaussians with excessively large scales or improper opacity values. To address this issue, we propose applying a 3DGS filter to exclude these floaters before the stylization process. After filtering certain 3D Gaussians, we fine-tune $G_{\theta}^{rec}$ to minimize the impact on the overall scene reconstruction quality and enhance color correspondence. During fine-tuning, we incorporate the reconstruction loss from [17]:
$\mathcal{L}_{rec}=(1-\lambda_{rec})\mathcal{L}_{1}(\mathcal{I}_{content}^{re},\mathcal{I}_{render})+\lambda_{rec}\mathcal{L}_{D-SSIM}$ (2)
Stylization. After matching the color and filtering possible artifacts from 3DGS reconstruction, an intuitive and effective strategy for 3D scene stylization is to leverage features extracted by a pre-trained convolutional neural network (e.g., VGG [28]) and calculate a set of loss functions between training views and the style image. To transfer detailed high-frequency style features from a 2D style image to a 3D scene, we exploit the nearest neighbor feature matching concept inspired by [13, 43], where we minimize the cosine distance between the feature map of rendered images and its nearest neighbor in the style feature map. As illustrated in Fig. 2, we extract the VGG feature maps $F_{style}$, $F_{content}$, and $F_{render}$ for $\mathcal{I}_{style}$, $\mathcal{I}_{content}^{re}$, and the rendered images $\mathcal{I}_{render}$, respectively. The nearest neighbor feature match loss is formulated as:
$\mathcal{L}_{style}(F_{render},F_{style})=\frac{1}{N}\sum_{i,j}^{N}D(F_{render}(i,j),F_{style}(i*,j*))$ (3)
[…]
Here, $D(a,b)$ is the cosine distance between two feature vectors $a$ and $b$. To preserve the original content structure during the stylization, we additionally minimize the mean squared distance between the content feature $F_{content}$ and the stylized feature $F_{render}$:
$\mathcal{L}_{content}=\frac{1}{H\times W}||F_{content}-F_{render}||^{2}$ (5)
where $H$ and $W$ represent the size of rendered images. While the content loss $\mathcal{L}_{content}$ mitigates the risk of excessive stylization, optimizations applied to 3DGS geometric parameters can still induce changes in scene geometry. To address this, we introduce a depth preservation loss to maintain general geometric consistency without relying on an additional depth estimator. Therefore, there exists a trade-off between geometric optimization and depth preservation loss, enabling our stylization to effectively transfer style patterns while mitigating substantial detriment to geometric content. Specifically, in the color match stage, we initially generate the original depth map $D_{origin}$ by employing the alpha-blending method of 3DGS. The alpha-blended depth $d^{\alpha}$ of each pixel is calculated as follows:
$d^{\alpha}=\sum_{i}T_{i}\alpha_{i}d_{i}$ (6)
where $d_{i}$ is the depth value for $i$-th Gaussian. During the stylization, we minimize the $L_{2}$ loss between the rendered depth map $D_{render}$ and the original one $D_{origin}$:
$\mathcal{L}_{depth}=\frac{1}{H\times W}||D_{origin}-D_{render}||^{2}$ (7)
We also add some regularization terms that perform on the changes of scale $\Delta s$ and opacity $\Delta\alpha$ value in $G_{\theta}^{sty}$:
$\mathcal{L}_{reg}^{ds}=\frac{1}{M}||\Delta s||,\quad\mathcal{L}_{reg}^{d\alpha}=\frac{1}{M}||\Delta\alpha||$ (8)
Finally, a total variation loss $\mathcal{L}_{tv}$ is used to smooth the rendered images in the 2D domain. The total loss during the stylization phase is:
$\mathcal{L}=\lambda_{sty}\mathcal{L}_{style}+\lambda_{con}\mathcal{L}_{content}+\lambda_{dep}\mathcal{L}_{depth}$ (9)
$+\lambda_{sca}\mathcal{L}_{reg}^{ds}+\lambda_{opa}\mathcal{L}_{reg}^{d\alpha}+\lambda_{tv}\mathcal{L}_{tv}$
where $\lambda_{*}$ is the corresponding loss weight.
Figure 3: Color control Results. Our approach facilitates versatile color management in stylized outputs, allowing users to either retain the scene’s original hues or apply distinct color schemes from alternative style images. Users have the autonomy to transfer the entire style, the pattern style only, or mixed arbitrary pattern and color styles.
3.3 Perceptual Control
With the previous components of our approach, we can achieve both rapid and detailed style transfer while preserving the original content. Here, we attempt to address the challenge users encounter in customizing the stylization process. Drawing inspiration from [22], we propose a series of strategies and loss functions for 3DGS stylization to control various perceptual factors, including color, scale, and spatial areas, aiming to enhance the customization capability of our approach.
Color Control. The color information within an image is a crucial perceptual aspect of its style. Unlike other stylistic elements such as brush strokes or dominant geometric shapes, color is largely independent and stands out as a distinct characteristic. As illustrated in Fig. 3, users may seek to preserve the original color scheme by only transferring the pattern style, or combine the pattern style of one style image with the color style of another. Consequently, our method provides users with independent control over color style during the stylization process. To achieve this, in our color transfer step, we can pre-color 3D scenes to align with any user-desired hues. Similar to the luminance-only algorithm proposed in [44], we then utilize the YIQ color space to separate the learning of luminance and color information. As shown in the definition of color style loss $\mathcal{L}_{style}^{color}$ in Eq. 10, we extract the luminance channel in the rendered view $\mathcal{I}_{render}^{Y}$ and the style image $\mathcal{I}_{style}^{Y}$ for the luminance-only style loss $L_{style}^{lum}$ calculation. Meanwhile, the RGB channels are retained for the content loss $\mathcal{L}_{content}$ calculation.
$\mathcal{L}_{style}^{color}=\mathcal{L}_{style}(F(\mathcal{I}_{render}^{Y}),F(\mathcal{I}_{style}^{Y}))$ (10)
Figure 4: Scale Control Results. Our method supports users to flexibly control the scale of the basic style elements, such as increasing or decreasing the density of Lego blocks in the example of last row.
Scale Control. The style scale, such as the thickness of brushstrokes or the density of the basic grains (as depicted in Fig. 4), is a foundational style element and plays a vital role in defining visual aesthetics. The stroke size, as a typical example of the style scale, is influenced by two factors: the receptive field of the convolution in the VGG network and the size of the style image input to the VGG network, as identified by [23]. Our exploration reveals that simply adjusting the receptive field is efficient for achieving scale-controllable stylization in 3DGS. We modulate the size of the receptive field by manipulating the selection of layers in the VGG network and adjusting the weights for each layer during the computation of the scale style loss $\mathcal{L}_{style}^{scale}$. As shown in Eq. 11, the style losses from different layer blocks are amalgamated to form $\mathcal{L}_{style}^{scale}$.
$\mathcal{L}_{style}^{scale}=\sum_{l\in b_{s}}w_{l}\cdot\mathcal{L}_{style}(F_{l}(\mathcal{I}_{render}),F_{l}(\mathcal{I}_{style}))$ (11)
Here, $F_{l}$ represents the feature map at the $l$-th layer within the $s$-th VGG-16 block $b_{s}$, and $w_{l}$ controls the corresponding weights at the $l$-th layer.
Figure 5: Spatial Control Results. By specifying regions in a single view and the corresponding style images for different regions, users can transfer different styles to desired regions within a scene. The users can specify some points (in blue) to generate region masks (in black).
Spatial Control. Traditional 3D style transfer methods typically apply a single style uniformly across the entire scene, resulting in the same stylization across the composition. However, there are scenarios where it becomes necessary to transfer distinct styles to different areas or objects. Users may also want to specify region-to-region constraints for their preferred style distribution within a scene, as illustrated in Fig. 5. Users only need to specify some points in the desired region of a single view and then our method utilizes SAM [45] to extract masks with those specified points to ease the interaction burden. Similarly, users can also specify certain areas in the style image to delineate style regions for partial style transfer. Additionally, we introduce a mask-tracking strategy that dynamically adjusts the position of input points based on area thresholds, ensuring the consistent generation of masks across various training views. More details can be found in the supplementary document.
To provide spatial guidance during the stylization process, we introduce a spatial style loss, denoted as $\mathcal{L}_{style}^{spatial}$. Assuming the user specifies $r$ regions in a single view of a scene to be matched with $r$ style regions, our $\mathcal{L}_{style}^{spatial}$ is formulated as follows:
$\mathcal{L}_{style}^{spatial}=\sum_{r}w_{r}\cdot\mathcal{L}_{style}(F_{b_{s}}(M_{r}^{c}\circ\mathcal{I}_{render}),F_{b_{s}}(M_{r}^{s}\circ\mathcal{I}_{style}^{r}))$ (12)
Here, $M_{r}^{c}$ and $M_{r}^{s}$ represent the binary spatial masks on the rendered views and the $r$-th style image, respectively. Additionally, during the color match step, we transform the color information in the region of the style image to the corresponding masked area of the content image to improve color correspondence.
Figure 6: Sequential control results. Given multiple control conditions, we can achieve a sequence of controllable stylization. We first show a stylization result and then, from left to right, progressively implement controls that preserve the color of the original scene, increase the scale of the pattern style, and adopt spatial control to apply two styles.
Figure 7: Qualitative comparisons with the baseline methods on LLFF and T&T datasets. It can be seen that our method has better results which faithfully capture both the color styles and pattern styles across different views.
Figure 8: Qualitative comparisons with baselines on LLFF dataset. Compared to other methods, our method excels in learning intricate and accurate geometric patterns while effectively preserving the semantic content of the original scene.
4 Experiments
Datasets. We conduct extensive experiments on multiple real-world scenes from LLFF [46] which contains forward-facing scenes and Tanks & Temples (T&T) dataset [47] which includes unbounded ${}^{\circ}$ large outdoor scenes. The scene data is also employed to train photo-realistic 3DGS representation before stylization with official configuration. Furthermore, our experiments involve a diverse set of style images from [13], allowing us to assess our method’s capability to handle a wide range of stylistic exemplars.
Baseline. We compare our method with the state-of-the-art 3D stylization methods, ARF [13], Stylescene [7], StyleRF [48] and Ref-NPR [42] based on different 3D representations. Specifically, Stylescene [7] adopts point clouds featured by VGG features averaged across views as a scene representation and transforms the point-wise features by modulating them with the encoding vector of a style image for stylization. ARF [13] is a NeRF-based approach that performs nearest neighbor feature matching on VGG feature maps along with deferred back propagation to memorize efficiently optimized parameters. StyleRF [48] implements style transfer within the feature space of radiance fields in a zero-shot manner. Ref-NPR [42] facilitates controllable scene stylization utilizing radiance fields to stylize a 3D scene, with a single stylized 2D view taken as reference. For all methods, we use their released code and pre-trained models. As Ref-NPR is a reference-based method, we use AdaIN [49] to obtain a stylized reference view according to its paper.
Implementation Details. We implement StylizedGS based on 3DGS representation [17]. In the color match step, we fix the density component of the initial 3DGS and optimize it using $\mathcal{L}_{rec}$ for 200 iterations. During the stylization process, we fine-tune all parameters of $G_{\theta}^{sty}$ over 800 iterations, while disabling adaptive density control of 3DGS. We adopt a pretrained VGG-16 network [28] to extract feature maps $F_{render}$, $F_{style}$, and $F_{content}$ used in Eq. 3. The network consists of 5 convolution blocks and we use the ‘conv2’ and ‘conv3’ blocks as the feature extractors empirically.
The loss weights $(\lambda_{rec},\lambda_{sty},\lambda_{con},\lambda_{sca},\lambda_{opa},\lambda_{tv})$ are respectively set to $(0.2,2,0.005,0.05,0.05,0.02)$. The depth preservation loss weight $\lambda_{dep}$ is set to $0.01$ for all forward-facing captures and $0.05$ for all ${}^{\circ}$ captures. At each stylization training iteration, we render an image from a sampled view different from all the training views to compute losses. The Adam optimizer [50] is adopted with a learning rate exponentially decayed from $11e-$ 1 to $21e-$ 2.
4.1 Qualitative Comparisons
We show visual comparisons with Stylescene and ARF methods in Fig. 7. We can see that our method exhibits a better style match to the style image compared to the others. For instance, Stylescene tends to generate over-smoothed results and lacks intricate structures in the Truck scene, while ARF struggles with propagating the style consistently across the entire scene and exhibits deficiencies in color correspondence. In contrast, our method faithfully captures both the color tones and brushstrokes of “The Scream” throughout the entire scene, demonstrating a superior style matching compared to the baselines.
Fig. 8 presents additional comparison results between ARF, StyleRF, Ref-NPR, and our method on LLFF datasets. By jointly optimizing the density and color components, our method excels in learning intricate geometric patterns and meso-structure. Specifically, in the orchids scene with a Lego style, other methods captures only the color style or a basic square pattern, whereas our approach successfully illustrates the detailed structure of Lego blocks, clearly showing even the small cylindrical patterns. This is also evident in the horns scene featuring a snow style, where our method produces more visually pleasant results that faithfully match the style of the hexagonal snowflake structure. Our method also preserves more semantic information in the stylized scene as we utilize the depth map to keep the spatial distribution in an image.
TABLE I: Quantitative comparisons on stylization under novel views. We report ArtFID, SSIM, average training time (Avg. train), and average rendering FPS (Avg. FPS) for our method and other baselines. ‘-’ denotes StyleRF does not require individual stylization training.
Metrics ArtFID($\downarrow$) SSIM($\uparrow$) Avg. train($\downarrow$) Avg. FPS($\uparrow$)
Stylescene 52.75 0.13 6.37 min 0.71
StyleRF 40.61 0.41 - 5.8
ARF 35.73 0.29 1.78 min 8.2
Ref-NPR 33.56 0.31 2.31 min 7.3
Ours 28.29 0.55 0.87 min 153
4.2 Quantitative Comparisons
Another significant advantage of the 3DGS representation is that it can produce high-quality renderings in real time and greatly reduce training time. Therefore, we conducted quantitative comparisons with other methods to demonstrate both the high fidelity and efficiency of our approach. As the optimization time during the stylization process is notably influenced by the resolution, we randomly selected 50 stylized scenes with consistent resolution from the LLFF dataset and evaluated the average times. For the stylization metrics, we choose ArtFID [51] to measure the quality of stylization following [52] and SSIM (Structural Similarity Index Measure) between the content and stylized images to measure the performance of detail preservation following [33]. As shown in Tab. I, our method excels in both stylized visual quality and content preservation. Regarding efficiency, our method has significant advantages in average rendering FPS (Avg. FPS) and can achieve real-time free-view synthesis. At the same time, it also outperforms other optimization-based stylization methods in terms of average training time (Avg. train) for stylization and can achieve per scene stylization within 1 minute on a single NVIDIA RTX 4090 GPU. Note that StyleRF is a zero-shot method and does not require individual stylization training for each case, but its rendering efficiency is worse and cannot be viewed in real time.
Figure 9: User Study. We record the user preference in the form of a boxplot. Our results obtain more preferences in visual preference, style match level, content preservation level, and 3D consistency quality than other state-of-the-art stylization methods.
User Study We also perform a user study among our method and other comparison methods. We randomly chose 20 sets of stylized views of the 3D scenes from both the LLFF and T&T datasets processed by different methods and invited 25 participants (including 15 males and 10 females, aged from 18 to 50). Our questionnaire was thoughtfully designed to encompass diverse participants across various professions and age groups, minimizing the risk of echo chamber effects. Participants were granted ample time for unbiased evaluation, ensuring a comprehensive and fair assessment. First, we presented the participants with a style image, the original scene, as well as two stylized videos generated by our method and a random comparison method. Then we asked the participants to vote on the video based on four evaluation indicators: visual preference, style match level, content preservation level, and 3D consistency quality. We collected 1000 votes for all evaluating indicators and presented the result in Fig. 9 in the form of boxplot. Our method, StylizedGS, outperforms those of other methods across all evaluation indicators, resulting in visually preferable and consistent results. A p-value of 0.005 indicates the statistical significance of the obtained results.
Figure 10: Ablation Study about density and color control. ‘w/o density’ presents the stylized scene without density component fine-tuning in 3DGS. ‘w/o recolor’ displays the stylized results without applying recoloring while ‘w/o depth’ shows the results without incorporating depth loss, leading to the disappearance of original geometry and semantic content.
Figure 11: Ablation study about 3D Gaussian filter. The filter effectively helps eliminate artifacts.
4.3 Controllable Stylization Results
Our method enables users to control stylization in three ways, including color, scale, and spatial region. The corresponding controllable stylization results are shown in Figs. 3, 4 and 5, respectively. Fig. 6 further demonstrates our method’s sequential control capability, enabling precise and flexible adjustments across different conceptual factors. In addition to the direct stylization result, we apply different controls sequentially from left to right: preserving the color of the original scene, increasing the scale of the pattern style, and adopting spatial control to apply two styles. This flexibility allows for refined modifications of key parameters or stylistic attributes, tailored to specific aesthetic preferences or artistic requirements. This empowers users to achieve finely tailored adjustments, fostering enhanced artistic expression and creative exploration.
TABLE II: Quantitative ablation study results. We report ArtFID and SSIM for our method and other ablation studies over 50 randomly chosen stylized cases.
Metrics w/o recolor w/o density w/o depth w/o GS filter Ours
ArtFID ($\downarrow$) 31.57 38.54 31.20 30.02 28.29
SSIM ($\uparrow$) 0.54 0.53 0.35 0.39 0.55
4.4 Ablation Study
We perform ablation studies to validate our design choices. The original scene and input style image are presented in Fig. 10 (a). Without applying the recoloring procedure, the stylized scene in (b) shows inferior color correspondence with the style image. In Fig. 10 (d)(e), we illustrate the importance of our depth preservation loss in preserving the original scene’s geometry. Without applying depth loss, the background of the Horse scene disappears, and the Trex’s skeleton becomes blurrier. The heatmap in the top left corner depicts the difference between the rendered depth maps and ground truth depth maps. The significant disparities in the heatmap of (d) also underscore the effectiveness of our depth preservation loss. In Fig. 10 (c), we highlight the importance of fine-tuning both the color and density components in 3DGS, which is crucial for learning intricate style patterns. Neglecting the optimization of the density makes it challenging to capture stroke details in the sketch and shape patterns in the texture style. Therefore, a delicate equilibrium is established wherein fine-tuning the density component harmonizes with the depth preservation loss, thereby facilitating the transfer of intricate style details while safeguarding the integrity of the original geometry. The quantitative results of ablation studies are presented in Tab.  II which also validate the significant impact of each component on the performance of our proposed method.
Additionally, we explore the impact of the 3DGS filter, addressing potential floaters in the reconstructed scenes. The floaters become more noticeable after the style transfer process. To mitigate this, we introduce a pre-filtering process. Specifically, we filter out 3D Gaussians with opacity below a certain threshold before we perform the style transfer fine-tuning process. Without the pre-filtering process, the empty space will be given textures. Fig. 11 shows the comparisons and we highlight some noticeable local areas with red bounding boxes. The first row presents results without pre-filtering, where the stone’s and leaves’ textures extend into the air around the objects. In the second row, after pre-filtering and training, the scene exhibits enhanced clarity, with fewer visible colored floaters. Tab. II shows the quantitative results of the ablation studies which demonstrate that all components contribute to our proposed method.
5 Discussion and Conclusion
Limitations. Although our method achieves efficient and controllable 3DGS stylization, it still has some limitations. First, the geometric artifacts from the original 3DGS reconstruction may impact the quality of the final stylized scenes. Although our 3DGS filter can eliminate some floaters, it cannot eliminate all geometric artifacts. This can rely on some improved 3DGS reconstruction methods that add geometric constraints, such as SuGaR [53]. Another point is that our optimization-based method cannot achieve instant stylization. However, this does not compromise our method’s stylization effect and practical application as the optimization time is controlled at about 1 minute.
In this paper, we propose StylizedGS, the first controllable 3D scene stylization method based on 3D Gaussian Splatting representation. After the 3DGS reconstruction, the proposed GS filter minimizes the impact of floaters in the reconstruction, which is crucial for the final stylization effect. Then we propose to adopt the nearest-neighbor feature matching style loss to optimize both geometry and color parameters of 3D Gaussians to capture the detailed style feature and facilitate the stylization of the 3D scene. We further introduce a depth preservation loss for regularization to preserve the overall structure during the stylization. Moreover, we design controllable ways for three perceptual factors: color, stylization scale, and spatial regions, so our method supports specific and diverse user control. Qualitative and quantitative experiments demonstrate that our method outperforms existing 3D stylization methods in terms of effectiveness and efficiency.